A tank is filled with 0.025 m3 of refrigerant 134-a refrigerant in the form of saturated liquid at 0.9 MPa on one side. The refrigerant then fills up the entire tank and it becomes 25 °C and 300 kPa. Calculate the total volume of the tank.


1
Expert's answer
2022-02-11T06:09:02-0500

Total volume

"V_2 =V_1(\frac{p_1}{p_2})^{\frac{1}{\lambda }}"

"V_2 =0.025(\frac{900}{300})^{\frac{1}{1.4}}= 0.055 m^3"